### v1

@@ Abstract
The randomized singular value decomposition (SVD) has become a popular approach to computing cheap, yet accurate, low-rank approximations to matrices due to its efficiency and strong theoretical guarantees. Recent work by Boullé and Townsend (FoCM, 2023) presents an infinite-dimensional analog of the randomized SVD to approximate Hilbert–Schmidt operators. However, many applications involve computing low-rank approximations to symmetric positive semi-definite matrices. In this setting, it is well-established that the randomized Nyström approximation is usually preferred over the randomized SVD. This paper explores an infinite-dimensional analog of the Nyström approximation to compute low-rank approximations to non-negative self-adjoint trace-class operators. We present an analysis of the method and, along the way, improve the existing infinite-dimensional bounds for the randomized SVD. Our analysis yields bounds on the expected value and tail bounds for the Nyström approximation error in the operator, trace, and Hilbert-Schmidt norms. Numerical experiments for simple integral operators validate the proposed framework.
@@ 1 Introduction
$\widehat{\bm{A}}:=\bm{A}\bm{\Omega}(\bm{\Omega}^{*}\bm{A}\bm{\Omega})^{\dagger}\bm{\Omega}^{*}\bm{A}\approx\bm{A},$ (1)
where $\bm{\Omega}$ is, again, a random sketch matrix and $\dagger$ denotes the Moore–Penrose pseudoinverse of a matrix. This approximation has received significant attention in the literature [19, 32, 39] and, like the randomized SVD, it enjoys strong theoretical guarantees. With the same number of matrix-vector products with $\bm{A}$, the randomized Nyström approximation is provably more efficient than the randomized SVD applied to an SPSD matrix; see e.g. [38, Lem. 5.2] and [33, Thms. 3.2 and 3.5].
Recently, Boullé and Townsend [7, 8] generalized the randomized SVD from matrices to Hilbert–Schmidt operators. Subsequent works [6, 10] employed this infinite-dimensional generalization of the randomized SVD to learn Green’s functions associated with an elliptic or parabolic partial differential equations (PDE) from a few solutions of the PDE. This approach uses hierarchical low-rank techniques and exploits the fact that Green’s functions are smooth away from the diagonal and therefore admit accurate off-diagonal low-rank approximations [3, 4]. The connection between numerical linear algebra and operator learning [9] motivates the development of infinite-dimensional analog of low-rank approximation algorithms, such as the randomized Nyström approximation, to derive sample complexity estimates of neural operator methods [24, 26, 27] in scientific machine learning. Other applications, like Gaussian process regression and Support Vector Machines [16, 18, 31, 40, 42, 43], involve integral operators that feature positive and globally smooth kernels. In turn, the operator is not only self-adjoint and positive but it also allows for directly applying low-rank approximation, without the need to resort to hierarchical techniques. Given existing results on matrices, it would be sensible to use an infinite-dimensional extension of the randomized Nyström approximation in such situations.
@@ 1.1 Contributions
Therefore, to analyze the infinite-dimensional generalization of the Nyström approximation, we will proceed through an analysis of the finite-dimensional Nyström approximation (1) when the columns of $\bm{\Omega}$ are drawn from a non-standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{K})$, for some general symmetric positive semi-definite matrix $\bm{K}$. We derive expectation and probability bounds between the matrix and its approximant in the Frobenius, spectral, and nuclear norms in Theorems 2.1, 2.4, and 2.6. Then, we use continuity arguments to obtain bounds on the infinite-dimensional generalization of Eq. 1 in Theorem 3.4. As a byproduct of our analysis, we improve the analysis of the infinite-dimensional analog of the randomized SVD. In particular, unlike the bounds presented in [8], our improved bounds coincide with the bounds obtained by Halko, Martinsson, and Tropp [20] in the finite-dimensional case when $\bm{K}$ is chosen as the identity matrix.
@@ 1.2 Organization of the paper
The paper is organized as follows. In Section 2 we present the setting for the Nyström approximation in finite dimensions, give an overview of Gaussian random vectors, and provide the analysis of the Nyström approximation for matrices with correlated Gaussian sketches. We then present the generalization of the Nyström approximation to Hilbert-Schmidt operators in Section 3. Next, we conclude with numerical experiments in Section 4.
@@ 2 Nyström approximation in finite dimensions with correlated Gaussian sketches
In this section, we present an analysis of the randomized Nyström approximation defined in (1), where the columns of $\bm{\Omega}$ are independent and identically distributed (i.i.d.) random vectors with distribution $\mathcal{N}(\bm{0},\bm{K})$ for some SPSD covariance matrix $\bm{K}$. We derive bounds in the Frobenius, spectral, and nuclear norms, which we denote respectively by $\|\bm{A}\|_{\textup{F}}=(\sum_{i=1}^{n}\sigma_{i}^{2})^{1/2}$, $\|\bm{A}\|_{2}=\sigma_{1}$, and $\|\bm{A}\|_{*}=\sum_{i=1}^{n}\sigma_{i}$, where $\sigma_{1}\geq\sigma_{2}\geq\ldots,\geq\sigma_{n}$ are the singular values of $\bm{A}$. More generally, $\|\bm{A}\|_{(s)}=(\sum_{i=1}^{n}\sigma_{i}^{s})^{1/s}$ denotes the Schatten $s$-norm, which coincides with the Frobenius norm when $s=2$ and the nuclear norm when $s=1$.
@@
$\bm{A}=\begin{bmatrix}\bm{U}_{1}&\bm{U}_{2}\end{bmatrix}\begin{bmatrix}\bm{\Sigma}_{1}&\bm{0}\\ \bm{0}&\bm{\Sigma}_{2}\end{bmatrix}\begin{bmatrix}\bm{U}_{1}^{*}\\ \bm{U}_{2}^{*}\end{bmatrix},\quad\bm{\Sigma}_{1}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{k}),\quad\bm{\Sigma}_{2}=\operatorname{diag}(\sigma_{k+1},\ldots,\sigma_{n}),$ (2)
@@ 2.1 Distribution of Gaussian sketches
$\widetilde{\bm{K}}=\begin{bmatrix}\bm{U}_{1}^{*}\bm{K}\bm{U}_{1}&\bm{U}_{1}^{*}\bm{K}\bm{U}_{2}\\ \bm{U}_{2}^{*}\bm{K}\bm{U}_{1}&\bm{U}_{2}^{*}\bm{K}\bm{U}_{2}\end{bmatrix}=:\begin{bmatrix}\widetilde{\bm{K}}_{11}&\widetilde{\bm{K}}_{21}^{*}\\ \widetilde{\bm{K}}_{21}&\widetilde{\bm{K}}_{22}\end{bmatrix},\quad\widetilde{\bm{K}}_{11}\in\mathbb{R}^{k\times k}.$ (3)
We assume that $\widetilde{\bm{K}}_{11}$ is invertible, which allows us to define the Schur complement $\widetilde{\bm{K}}_{22.1}=\widetilde{\bm{K}}_{22}-\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}$. By well-known properties of Gaussian random vectors [21, Thm. 5.2], $\bm{\omega}\sim\mathcal{N}(0,\bm{K})$ implies $\bm{U}^{*}\bm{\omega}\sim\mathcal{N}(0,\widetilde{\bm{K}})$ and, thus, the marginal distributions
@@
In particular, the columns of $\bm{\Omega}_{1}$ are i.i.d. as $N(\bm{0},\widetilde{\bm{K}}_{11})$ or, in other words, $\bm{\Omega}_{1}=\smash{\widetilde{\bm{K}}_{11}^{1/2}\bm{X}}$ with the standard Gaussian matrix $\bm{X}\in\mathbb{R}^{k\times(k+p)}$. Because $\widetilde{\bm{K}}_{11}$ is invertible, this shows that, with probability one, $\bm{\Omega}_{1}$ has full rank and possesses a right inverse $\bm{\Omega}_{1}^{\dagger}$, which satisfies $\bm{\Omega}_{1}\bm{\Omega}_{1}^{\dagger}=\bm{I}_{k}$. Finally, the conditional probability distribution of $\bm{\omega}_{2}$ given $\bm{\omega}_{1}=\bm{x}_{1}$ is normal with mean $\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{x}_{1}$ and covariance matrix $\widetilde{\bm{K}}_{22.1}$ [21, Thm. 5.3].
@@ 2.2 Nyström method with correlated Gaussian sketches
Similarly to the error analysis for the randomized SVD with correlated input vectors [7, 8], our error bounds for the Nyström approximation depend on prior information contained in the representation Eq. 3 of $\bm{K}$ with respect to the eigenvectors of $\bm{A}$. This is measured through the following two quantities:
$\beta_{k}^{(\xi)}=\frac{\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}\bm{\Sigma}_{2}^{1/2}\|_{\xi}}{\|\bm{\Sigma}_{2}\|_{\xi}}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2},\quad\delta_{k}^{(\xi)}=\frac{\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2}\|_{\xi}}{\|\bm{\Sigma}_{2}\|_{\xi}}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2},$ (4)
where $\xi\in\{\textup{F},2,*\}$ such that $\|\cdot\|_{\xi}$ dictates the choice of norm. The following theorem states our main result concerning the approximation error of the Nyström approximation in the Frobenius norm. Similar results are derived in the spectral and nuclear norms but are deferred to Section 2.2.3 for clarity of exposition.
Theorem 2.1 (Frobenius norm).
Let $\bm{A}$ be an $n\times n$ SPSD matrix, $2\leq k\leq\operatorname{rank}(\bm{A})$ be a target rank, and $p\geq 4$ an oversampling parameter. Let $\bm{\Omega}$ be a random sketch matrix with the columns i.i.d. $\mathcal{N}(\bm{0},\bm{K})$, where the covariance matrix $\bm{K}$ is such that the matrix $\widetilde{\bm{K}}_{11}$ defined in Eq. 3 is invertible. Then, the Nyström approximation $\widehat{\bm{A}}$ satisfies
$\mathbb{E}[\|\bm{A}-\widehat{\bm{A}}\|_{\textup{F}}]\leq\left(1+2\delta_{k}^{(\textup{F})}+2\sqrt{c_{1}}\beta_{k}^{(\textup{F})}\right)\|\bm{\Sigma}_{2}\|_{\textup{F}}+2\sqrt{c_{2}}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*},$ (5)
where $c_{1}=\mathcal{O}(k^{2}/p^{2})$, $c_{2}=\mathcal{O}(k^{2}/p^{2})$ are constants defined later in Eq. 9. Let $u,t\geq 1$, then
$\|\bm{A}-\widehat{\bm{A}}\|_{\textup{F}}$ $\leq\|\bm{\Sigma}_{2}\|_{\textup{F}}+4\left(\delta_{k}^{(\textup{F})}+t^{2}(d_{1}+d_{3})\beta_{k}^{(\textup{F})}\right)\|\bm{\Sigma}_{2}\|_{\textup{F}}+4t^{2}d_{3}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*}$ (6)
$\quad+2t^{2}u^{2}d_{2}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2}$
holds with probability at least $1-3t^{-p}-e^{-u^{2}/2}$. Here, $d_{1}=\mathcal{O}(k/p)$, $d_{2}=\mathcal{O}(k/p)$, and $d_{3}=\mathcal{O}(k^{3/2}/p)$ are constants defined in Eq. 12.
The proof of Theorem 2.1 is based on an existing structural bound:
$\|\bm{A}-\widehat{\bm{A}}\|\leq\|\bm{\Sigma}_{2}\|+\|(\bm{\Sigma}_{2}^{1/2}\bm{\bm{\Omega}}_{2}\bm{\Omega}_{1}^{\dagger})^{*}\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|,$ (7)
@@
2.2.1 Proof of the expectation bound
To prove inequality Eq. 5 of Theorem 2.1, we use the $L_{q}$ norm of a random variable $Z$ defined as $\mathbb{E}^{q}(Z)=\mathbb{E}[|Z|^{q}]^{1/q}$. After taking expectation with respect to $\bm{\Omega}$ and applying Hölder’s inequality, the second term in the bound Eq. 7 for the Frobenius norm satisfies
To proceed from here, we recall that Section 2.1 provides the conditional distribution of $\bm{\Omega}_{2}|\bm{\Omega}_{1}$ as $\bm{\Omega}_{2}|\bm{\Omega}_{1}\sim\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{\Omega}_{1}+\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}$, where $\bm{\Psi}$ is a standard Gaussian matrix. Using the triangle inequality for the $L_{4}$ norm and $\bm{\Omega}_{1}\bm{\Omega}_{1}^{\dagger}=\bm{I}_{k}$ (which holds with probability $1$), we obtain that
$\mathbb{E}^{4}[\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}]$ $=\mathbb{E}_{\bm{\Omega}_{1},\bm{\Psi}}^{4}[\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{\Omega}_{1}\bm{\Omega}_{1}^{\dagger}+\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{(4)}]$ (8)
To bound the second term, we first apply Eq. 32b,
and then take expectation with respect to $\bm{\Omega}_{1}$ using Lemma A.5:
$\mathbb{E}_{\bm{\Omega}_{1}}\big{[}\mathbb{E}_{\bm{\Psi}}\big{[}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}\mid\bm{\Omega}_{1}\big{]}\big{]}\leq c_{1}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{2}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{4}+c_{2}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{2}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}^{4},$
where
$c_{1}=k\frac{(p-1)(k+1)+2}{p(p-1)(p-3)},\quad\text{and}\quad c_{2}=k\frac{k+p-1}{p(p-1)(p-3)}.$ (9)
Inserting the inequality above into Eq. 8 and using $\|\widetilde{\bm{K}}_{11}^{-1/2}\|_{2}=\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{1/2}$ gives
$\quad\qquad\big{(}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1/2}\|_{(4)}+\big{(}c_{1}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{4}+c_{2}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}^{4}\big{)}^{1/4}\big{)}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{1/2},$
where we used $(a+b)^{2}\leq 2a^{2}+2b^{2}$ and the subadditivity of the square-root. This concludes the proof of°Eq. 5.
2.2.2 Proof of the tailbound
We begin the proof of the tailbound (6) of Theorem 2.1 by a concentration inequality on norms of shifted and rescaled Gaussian matrices.
Lemma 2.2.
holds for every $u\geq 1$.
Proof 2.3.
Given $h(\bm{X}):=\|\bm{B}+\bm{C}\bm{X}\bm{D}\|_{(s)}$, we have
where we used that the Frobenius norm is larger than any Schatten-$s$ norm for $s\geq 2$. In other words, $h$ is Lipschitz continuous with Lipschitz constant $\|\bm{C}\|_{2}\|\bm{D}\|_{2}$. This allows us to apply a concentration result for functions of Gaussian matrices [20, Prop. 10.3], which yields the statement of the lemma.
To obtain Eq. 6, it suffices to derive a tailbound for the term $\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}$ in the structural bound Eq. 7 (in the Frobenius norm). Using that $\bm{\Omega}_{2}|\bm{\Omega}_{1}\sim\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{\Omega}_{1}+\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}$ with a standard Gaussian matrix $\bm{\Psi}$, Lemma 2.2 yields the following tailbound for $\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}$ conditioned on $\bm{\Omega}_{1}$:
$\mathbb{P}\left\{\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\geq\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\mid\bm{\Omega}_{1}\big{]}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{2}\|\bm{\Omega}_{1}^{\dagger}\|_{2}u\mid\bm{\Omega}_{1}\right\}\leq e^{-u^{2}/2}.$ (10)
In analogy to (8), combining the triangular inequality for the Schatten-$4$ norm and Jensen’s inequality yields
$\leq\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{(4)}+\left(\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{4}\big{(}\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}+\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}\big{)}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}^{4}\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}\right)^{1/4}.$
$\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\mid\bm{\Omega}_{1}\big{]}^{2}$ $\leq 2\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{(4)}^{2}+2\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{2}(\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}+\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{2})$ (11)
with the constants
$d_{1}=\frac{3k}{p+1},\quad d_{2}=e^{2}\frac{k+p}{(p+1)^{2}},\quad d_{3}=\sqrt{k}d_{2}.$ (12)
Lemma A.7 implies the following bound for $\mathbb{P}(E_{t}^{c})$:
$\mathbb{P}(E_{t}^{c})\leq 2t^{-(p+1)}+t^{-p}\leq 3t^{-p}.$ (13)
Under the event $E_{t}$, Eq. 11 becomes
$\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\mid E_{t}\big{]}^{2}\leq 2\left(\delta_{k}^{(\textup{F})}+t^{2}(d_{1}+d_{3})\beta_{k}^{(\textup{F})}\right)\|\bm{\Sigma}_{2}\|_{\textup{F}}+2t^{2}d_{3}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*}.$ (14)
Conditioning Eq. 10 on $E_{t}$ and combining it with Eq. 14 yields
with probability $\geq 1-e^{-u^{2}/2}$. Similarly to the proof of [20, Thm. 10.8], by the union bound we remove the conditioning by a union bound using Eq. 13 and conclude the proof of Theorem 2.1.
2.2.3 Analysis in the spectral and nuclear norms
The structural bound Eq. 7 for the Nyström method applies to any unitarily invariant matrix norm. Using properties of Gaussian matrices from Appendix A, this allows us to extend the analysis performed in Sections 2.2.1 and 2.2.2 to the spectral and nuclear norms.

### v2

@@ Abstract
The randomized singular value decomposition (SVD) has become a popular approach to computing cheap, yet accurate, low-rank approximations to matrices due to its efficiency and strong theoretical guarantees. Recent work by Boullé and Townsend (FoCM, 2023) presents an infinite-dimensional analog of the randomized SVD to approximate Hilbert-Schmidt operators. However, many applications involve computing low-rank approximations to symmetric positive semi-definite matrices. In this setting, it is well-established that the randomized Nyström approximation is usually preferred over the randomized SVD. This paper explores an infinite-dimensional analog of the Nyström approximation to compute low-rank approximations to non-negative self-adjoint trace-class operators. We present an analysis of the method and, along the way, improve the existing infinite-dimensional bounds for the randomized SVD. Our analysis yields bounds on the expected value and tail bounds for the Nyström approximation error in the operator, trace, and Hilbert-Schmidt norms. Numerical experiments on integral operators arising from Gaussian process sampling and Bayesian inverse problems are used to validate the proposed infinite-dimensional Nyström algorithm.
@@ 1 Introduction
(1) $\widehat{\bm{A}}:=\bm{A}\bm{\Omega}(\bm{\Omega}^{*}\bm{A}\bm{\Omega})^{\dagger}\bm{\Omega}^{*}\bm{A}\approx\bm{A},$
where $\bm{\Omega}$ is, again, a random sketch matrix and $\dagger$ denotes the Moore-Penrose pseudoinverse of a matrix. This approximation has received significant attention in the literature [20, 34, 40] and, like the randomized SVD, it enjoys strong theoretical guarantees. With the same number of matrix-vector products, the randomized Nyström approximation is typically significantly more accurate than the randomized SVD when the matrix has fast decaying singular values. Additionally, the Nyström method requires only a single pass over the matrix, compared to two passes for the randomized SVD, enabling all matrix-vector products to be performed in parallel.
Recently, Boullé and Townsend [7, 8] generalized the randomized SVD from matrices to Hilbert-Schmidt operators. Subsequent works [6, 10] employed this infinite-dimensional generalization of the randomized SVD to learn Green’s functions associated with an elliptic or parabolic partial differential equations (PDE) from a few solutions of the PDE. This approach uses hierarchical low-rank techniques and exploits the fact that Green’s functions are smooth away from the diagonal and therefore admit accurate off-diagonal low-rank approximations [4, 5]. The connection between numerical linear algebra and operator learning [9] motivates the development of infinite-dimensional analog of low-rank approximation algorithms, such as the randomized Nyström approximation, to derive sample complexity estimates of neural operator methods [26, 28, 29] in scientific machine learning. Other applications, like Gaussian process regression and Support Vector Machines [17, 19, 33, 41, 43, 44], involve integral operators that feature positive and globally smooth kernels. In turn, the operator is not only self-adjoint and positive but it also allows for directly applying low-rank approximation, without the need to resort to hierarchical techniques. Given existing results on matrices, it would be sensible to use an infinite-dimensional extension of the randomized Nyström approximation in such situations.
@@ 1.1 Contributions
Therefore, to analyze the infinite-dimensional generalization of the Nyström approximation, we will proceed through an analysis of the finite-dimensional Nyström approximation (1) when the columns of $\bm{\Omega}$ are drawn from a non-standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{K})$, for some general symmetric positive semi-definite matrix $\bm{K}$. We derive expectation and probability bounds between the matrix and its approximant in the Frobenius, spectral, and nuclear norms in Theorems 2.8, 2.1, and 2.5. Then, we use continuity arguments to obtain bounds on the infinite-dimensional generalization of (1) in Theorem 3.5. As a byproduct of our analysis, we improve the analysis of the infinite-dimensional analog of the randomized SVD. In particular, unlike the bounds presented in [8], our improved bounds coincide with the bounds obtained by Halko, Martinsson, and Tropp [21] in the finite-dimensional case when $\bm{K}$ is chosen as the identity matrix.
A key benefit of the proposed framework is that once a low-rank approximation of the operator is computed, one can obtain a low-rank approximation to any discretization of the operator at low additional cost. Clearly, an actual numerical implementation of our Nyström approximation also requires discretization to realize the action of the operator $\mathcal{A}$ on a function. However, this action can be approximated adaptively and to high accuracy, say up to machine precision, so that the error is negligible.
@@ 1.2 Organization of the paper
The paper is organized as follows. In Section 2 we present the setting for the Nyström approximation in finite dimensions, give an overview of Gaussian random vectors, and provide the analysis of the Nyström approximation for matrices with correlated Gaussian sketches. We then present the generalization of the Nyström approximation to Hilbert-Schmidt operators in Section 3. Next, we conclude with numerical experiments in Section 4. Our analysis relies on properties of random matrices with Gaussian random entries. The statements and proofs of these properties are deferred to Appendix A.
@@ 2 Nyström approximation in finite dimensions with correlated Gaussian sketches
In this section, we present an analysis of the randomized Nyström approximation defined in (1), where the columns of $\bm{\Omega}$ are independent and identically distributed (i.i.d.) random vectors with distribution $\mathcal{N}(\bm{0},\bm{K})$ for some SPSD covariance matrix $\bm{K}$. We derive bounds in the Frobenius, spectral, and nuclear norms, which we denote respectively by $\|\bm{A}\|_{\textup{F}}=(\sum_{i=1}^{n}\sigma_{i}^{2})^{1/2}$, $\|\bm{A}\|_{2}=\sigma_{1}$, and $\|\bm{A}\|_{*}=\sum_{i=1}^{n}\sigma_{i}$, where $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{n}$ are the singular values of $\bm{A}$. More generally, $\|\bm{A}\|_{(s)}=(\sum_{i=1}^{n}\sigma_{i}^{s})^{1/s}$ denotes the Schatten $s$-norm, which coincides with the Frobenius norm when $s=2$, the nuclear norm when $s=1$, and the spectral norm when $s=\infty$.
@@
(2) $\bm{A}=\begin{bmatrix}\bm{U}_{1}&\bm{U}_{2}\end{bmatrix}\begin{bmatrix}\bm{\Sigma}_{1}&\bm{0}\\ \bm{0}&\bm{\Sigma}_{2}\end{bmatrix}\begin{bmatrix}\bm{U}_{1}^{*}\\ \bm{U}_{2}^{*}\end{bmatrix},\quad\bm{\Sigma}_{1}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{k}),\quad\bm{\Sigma}_{2}=\operatorname{diag}(\sigma_{k+1},\ldots,\sigma_{n}),$
@@ 2.1 Distribution of Gaussian sketches
(3) $\widetilde{\bm{K}}=\begin{bmatrix}\bm{U}_{1}^{*}\bm{K}\bm{U}_{1}&\bm{U}_{1}^{*}\bm{K}\bm{U}_{2}\\ \bm{U}_{2}^{*}\bm{K}\bm{U}_{1}&\bm{U}_{2}^{*}\bm{K}\bm{U}_{2}\end{bmatrix}=:\begin{bmatrix}\widetilde{\bm{K}}_{11}&\widetilde{\bm{K}}_{21}^{*}\\ \widetilde{\bm{K}}_{21}&\widetilde{\bm{K}}_{22}\end{bmatrix},\quad\widetilde{\bm{K}}_{11}\in\mathbb{R}^{k\times k}.$
We assume that $\widetilde{\bm{K}}_{11}$ is invertible, which allows us to define the Schur complement $\widetilde{\bm{K}}_{22.1}=\widetilde{\bm{K}}_{22}-\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}$. By well-known properties of Gaussian random vectors [22, Thm. 5.2], $\bm{\omega}\sim\mathcal{N}(0,\bm{K})$ implies $\bm{U}^{*}\bm{\omega}\sim\mathcal{N}(0,\widetilde{\bm{K}})$ and, thus, the marginal distributions
@@
In particular, the columns of $\bm{\Omega}_{1}$ are i.i.d. as $N(\bm{0},\widetilde{\bm{K}}_{11})$ or, in other words, $\bm{\Omega}_{1}{\color[rgb]{0,0,0}\sim}\smash{\widetilde{\bm{K}}_{11}^{1/2}\bm{X}}$ with the standard Gaussian matrix $\bm{X}\in\mathbb{R}^{k\times(k+p)}$. Because $\widetilde{\bm{K}}_{11}$ is invertible, this shows that, with probability one, $\bm{\Omega}_{1}$ has full rank and possesses a right inverse $\bm{\Omega}_{1}^{\dagger}$, which satisfies $\bm{\Omega}_{1}\bm{\Omega}_{1}^{\dagger}=\bm{I}_{k}$. Finally, the conditional probability distribution of $\bm{\omega}_{2}$ given $\bm{\omega}_{1}=\bm{x}_{1}$ is normal with mean $\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{x}_{1}$ and covariance matrix $\widetilde{\bm{K}}_{22.1}$ [22, Thm. 5.3].
@@ 2.2 Nyström method with correlated Gaussian sketches
Similarly to the error analysis for the randomized SVD with correlated input vectors [7, 8], our error bounds for the Nyström approximation depend on prior information contained in the representation (3) of $\bm{K}$ with respect to the eigenvectors of $\bm{A}$. This is measured through the following two quantities:
(4) $\beta_{k}^{(\xi)}=\frac{\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}\bm{\Sigma}_{2}^{1/2}\|_{\xi}}{\|\bm{\Sigma}_{2}\|_{\xi}}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2},\quad{\color[rgb]{0,0,0}\delta_{k}^{(\xi)}=\frac{\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-2}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2}\|_{\xi}}{\|\bm{\Sigma}_{2}\|_{\xi}},}$
where $\xi\in\{\textup{F},2,*\}$ such that $\|\cdot\|_{\xi}$ dictates the choice of norm.
Note that $\widetilde{\bm{K}}_{22.1}$ is the covariance matrix of $\bm{\Omega}_{2}$ conditioned on $\bm{\Omega}_{1}$. Therefore, $\beta_{k}^{(\xi)}$ is small if the random vectors in $\bm{\Omega}$ have a small variance in the direction of the bottom eigenvectors of $\bm{A}$, conditioned on the randomness in the direction of the top $k$ eigenvectors of $\bm{A}$. In other words, $\beta_{k}^{(\xi)}$ measures the randomness in the eigenvectors corresponding to the lower eigenvalues of $\bm{A}$, weighted by the bottom $n-k$ eigenvalues of $\bm{A}$. Furthermore, as discussed in [13], note that $\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}=\tan(\bm{U}_{1},\bm{K}\bm{U}_{1})$ is a matrix whose singular values are the tangents of the principal angles between $\operatorname{range}(\bm{U}_{1})$ and $\operatorname{range}(\bm{K}\bm{U}_{1})$ [46, Theorem 3.1]. Therefore, to make $\delta_{k}^{(\xi)}$ small we should choose $\bm{K}$ so that $\bm{U}_{1}$ is an approximate invariance subspace for $\bm{K}$. Hence, $\delta_{k}^{(\xi)}$ is a measure of how well $\operatorname{range}(\bm{K}\bm{U}_{1})$ and $\operatorname{range}(\bm{U}_{1})$ align, weighted with the bottom $n-k$ eigenvalues of $\bm{A}$.
The rest of the section is devoted to proving error bounds for the Nyström approximation. The proofs of our error bounds are based on the existing structural bound
(5) $\|\bm{A}-\widehat{\bm{A}}\|\leq\|\bm{\Sigma}_{2}\|+\|(\bm{\Sigma}_{2}^{1/2}\bm{\bm{\Omega}}_{2}\bm{\Omega}_{1}^{\dagger})^{*}\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|,$
@@
2.3 Analysis in the nuclear and operator norms
The following theorem states our main result concerning the approximation error of the Nyström approximation in the nuclear and operator norms. Similar results are derived in the Frobenius norm but are deferred to Section 2.4 for clarity of exposition.
2.3.1 Expectation bounds
We begin with with the statement and proof of an expectation bounds for the error $\|\bm{A}-\widehat{\bm{A}}\|$ in the operator and nuclear norms.